x + y = 12
y = 2x
HELP

Answer by oberobic(2302)

One of the way to solve linear equations is to substitute.
In this case, you're given y=2x, so it's ready for substitution.
x+y=12 is the other equation, so substitute the value of y.
x+2x=12
Collecting like terms:
3x = 12
Dividing through by 3:
x = 4
Now go back to first equation: x + y = 12.
We now know x = 4, so it becomes:
4 + y = 12
Subtracting 4 from both sides:
y = 8.
ALWAYS check.
In this case, does y = 2x?
y = 2x
8 = 2*4
8 = 8.
Check!